The length of time it takes Brad to get to work varies inversely with his speed of travel. If it takes him 2.5 hours when traveling 15 miles per hour, how long does it take when traveling 50 miles per hour?
Answer by josmiceli(6786)

speed varies inversely with time
s = speed
t = time
k = proportional constant
s = k(1/t)
15 = k(1/2.5)
k = 2.5 * 15
k = 37.5
50 = 37.5(1/t)
t = 37.5/50
t = .75 or three quarters of an hour